OpenStudy (anonymous):

The equation below represents Function A and the graph represents Function B: Function A f(x) = x - 9 Function B graph of line going through ordered pairs negative 1, negative 3 and 2, 3 Which equation best compares the slopes of the two functions? Slope of Function B = 2 × Slope of Function A. Slope of Function A = Slope of Function B Slope of Function A = 2 × Slope of Function B Slope of Function B = - Slope of Function A

OpenStudy (anonymous):

@ganeshie8 @Leader @nincompoop @Nnesha

OpenStudy (leader):

Do you know the slope formula?

OpenStudy (anonymous):

y=mx+y @Leader

OpenStudy (leader):

\[\Large Slope= \frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (leader):

y=mx+b is the slope intercept form.

OpenStudy (anonymous):

oh

OpenStudy (leader):

And we are given y=x-9 so y=mx+b and m is the slope y=x-9 What is the slope of this one first? @heymydearpeople

OpenStudy (anonymous):

how do i find the slope

OpenStudy (anonymous):

@Leader

OpenStudy (leader):

y=x-9 and m is the slope of y=mx+b what number is multiplying x? Remember anything times 1 is itself :P

OpenStudy (anonymous):

y

OpenStudy (anonymous):

a or b is the answer right ?

OpenStudy (anonymous):

@Leader